Two records and three tapes cost $31. Three records and two tapes cost $29. Find the cost of each record and each tape.
Answer by checkley77(12844)

2R+3T=31 MUKTIPLY THIS EQUATION BY 3.
3R+2T=28 MULTIPLY THIS EQUATION BY -2 & ADD THEM.
6R+9T=93
-6R-4T=-56
---------------
5T=37
T=37/5
T=$7.40 FOR EACH TAPE.
2R+3*7.40=31
2R+22.20=31
2R=31-22.20
2R=8.8
R=8.8/2
R=$4.40 FOR EACH RECORD.
PROOF:
3*4.40+2*7.40=28
13.20+14.80=28
28=28